Find the disk of convergence of power series IM: (= –2+i) 2 where, z = x + iy n=0
Answer #1

(2-2 + ) ; where f/2) = Power Series - 2 zxriy nzo دوو (2-(2-2)) an nro 00 an nzo General form of buser dories - 8/2) = E ,we know whas modulus value is 2. {:: 12.- (2-1) ) = 2] I el = 1 ; Put z- (2-1) = 2 eo in in f(2) Imagine the point graphicall